A local hamburger shop sold a combined total of 626 hamburgers and cheeseburgers on Monday. There were 74 fewer cheeseburgers sold than hamburgers. How many hamburgers were sold on Monday?


Answer by jorel1380(3719)

h+c=626
h=c+74
c+74+c=626
2c=552
c=276
h=350
There were 350 hamburgers sold, and 276 cheeseburgers..